A 2 kilo-ohm resistor is connected in series to an 6 milli-henrie inductor and a 2000 V battery (DC). After, 1.22 micro-seconds, what is the voltage across the inductor in volts?
Answer #1

current in the circuit is given by

i = ( V/R) ( 1 - e ( - t R/ L))

i = (2000/ 2000) ( 1 - e^( - 1.22* 2000 * 10^-6/ ( 6* 10^-3))

i = 0.334 A

voltage across resistor

V = i R = 0.334* 2000 = 668.27 V

=========

Comment in case any doubt, will reply for sure.. Goodluck